3 equal point charges which are intially infinite apart are placed at the corners of an equallatiral triangle with sides d m long.thWhat is the elecntial tric potential energy of the system ?e two of the point charges are identical with charge q.If the zero net work is required to place the 3 charges at the cornerns then what would be the third charge ?

3 equal  point charges  which are intially infinite apart are placed at the corners of an equallatiral triangle with sides d m long.thWhat is the elecntial tric potential energy of the system ?e two of the point charges are identical with charge q.If the zero net work is required to place the 3 charges at the cornerns then what would be the third charge ?

Grade:12

1 Answers

vikas askiitian expert
13 years ago

let the vertices of triangle are a,b,c ...

ans for part 1)

total electric potential energy=kq2 /d +kq2 /d+kq2/d        (sum of potential energy of each pair)

                               (PE)total       =3kq2 /d

ans for part 2)

when two charges are idential & let other is Q..

now total potential energy=KQq/d +KQq/d+Kq2 /d           (sum of potential energy of each pair of charges)

        work required to place these charges at corner of equilateral triangle

         is equal to potential energy of system

so 

             total potential energy=total work=0

                    2kQq/d=-kq2/d

                       Q=-q/2 ans